# SCAWaveNet: A Spatial-Channel Attention-based Network for Global Significant Wave Height Retrieval

### v1

@@ III-B Lightweight Attention-based Auxiliary Parameters Branch
Then $A_{e}$ is split into two equal-sized vectors, denoted as $A_{1}$ and $A_{2}$, which are fed into two distinct attention blocks. Each block includes two projection layers and a sigmoid activation function, designed to extract attention weights along the spatial and channel dimensions. Specifically, $A_{1}$ undergoes an up-projection to a higher-dimensional space first, then is down-projected back to its original dimension, and finally passes through a sigmoid function to generate spatial attention weights. For $A_{2}$, it is transposed to feed into the attention block, and then transposed back to its original shape after weight extraction. This process is defined as:
@@
where $\otimes$ denotes element-wise multiplication. The output $A^{\prime}$ serves as the enhanced APs feature, integrating both spatial and channel-level characteristics.
@@ III-C Feature Fusion and Task Head
In the previous branches, features from the DDMs and APs were independently extracted to capture spatial and channel-wise information. These two types of features are concatenated along the channel dimension to form a unified representation. This feature is passed into the task head, which is an MLP with 9 hidden layers. The output includes SWH predictions across four channels, expressed as:
@@
where $y$ and $\hat{y}$ denote the ground truth and the predicted values, respectively, and $\delta$ is a hyperparameter that controls the transition between the Mean Squared Error (MSE) and Mean Absolute Error (MAE) characteristics. The Huber loss integrates the strengths of MSE and MAE: it is quadratic for small errors (ensuring smooth gradients) and linear for large errors (reducing outlier sensitivity). By tuning $\delta$, the loss function balances emphasis on large and small errors. In this study, the training objective is to optimize SWH predictions across all channels. Therefore, the loss is computed using the $\hat{y}$ and $y$ values of the four channels.
@@ IV-A Data Sources
In this study, GNSS-R observations are obtained from the CYGNSS Level 1 Version 3.2 science data product, provided by the Physical Oceanography Distributed Active Archive Center. CYGNSS, a NASA Earth System Science Pathfinder mission, consists of a constellation of eight microsatellites. These satellites provide near-continuous global coverage through an orbital inclination of approximately $35^{\circ}$ relative to the equator, with an average revisit time of 7 hours. This inclination enables CYGNSS to measure ocean parameters between nearly $38^{\circ}$ and $38^{\circ}$ latitude [https://doi.org/10.5067/cygns-l1x32].
One of the reference datasets used in this study is the ERA5 dataset, the fifth-generation global climate and weather reanalysis product developed by the European Centre for Medium-Range Weather Forecasts[45]. ERA5 provides three types of SWH products: the combined SWH of wind waves and swell, the SWH of wind waves alone, and the SWH of swell alone. In this study, the combined SWH data are employed, which offer a spatial resolution of $0.5^{\circ}\times 0.5^{\circ}$ and an hourly temporal resolution. These data serve as the foundation for model training, validation, and testing.
In parallel, buoy data obtained from the National Data Buoy Center (NDBC) are used as real-world observations. The NDBC operates 1,350 stations, of which 1,056 have reported in the past 8 hours. We specifically utilized the standard meteorological NDBC data, with each station reporting SWH measurements at 30- or 60-minute intervals. Due to the limited availability of buoy data, these measurements are used exclusively as a test set to evaluate the real-world performance of the SWH retrieval model.
@@ IV-B Data Preprocessing
After acquiring CYGNSS, ERA5 and NDBC data, a three-step data preprocessing procedure is applied: (1) quality control; (2) channel alignment; and (3) CYGNSS-ERA5 and CYGNSS-Buoy data matching.
@@ IV-B1 Quality Control
  • Remove observations where any of the first 28 bits in the quality control flag equals 1.
Figure 5: Data processing workflow used in this study.
@@ IV-B2 Channel Alignment
  • If data from any channel are missing at a timestamp, the entire observation for that timestamp is discarded.
@@ IV-B3 Data Matching
  • CYGNSS-Buoy data matching: The temporal and spatial matching thresholds for CYGNSS and NDBC buoy data are set to 30 minutes and 25 km, whereby the CYGNSS specular point must be within 25 km of the buoy location, and the observation time must fall within ± 30 minutes of the buoy measurement. Due to the stringent requirement of aligning all four CYGNSS channels to identical timestamps, data matching is performed independently for each channel.
Because the CYGNSS data exhibit signal saturation when SWH >8 m, these measurements offer limited modeling value. Therefore, we exclude all data points with SWH values above this threshold.
@@ IV-C Data Split
In this study, the temporal coverage of CYGNSS and ERA5 data spans from August 1, 2019, to July 31, 2022. To ensure fairness in model training and account for potential seasonal variations, the training, validation, and test sets were strictly divided into non-overlapping time periods. Specifically:
@@
  • The test set spans from August 1, 2021, to July 31, 2022.
Given that the dataset contains four-channel data, training a model on the full yearly dataset would be computationally expensive. Therefore, we randomly sampled 3 million, 0.5 million, and 3 million data points to form the final training, validation, and test sets. Furthermore, to ensure consistency in model evaluation, the buoy data are aligned with the temporal range of the test set.
@@ V-A Experimental Setup
  • CC quantifies the linear relationship between predictions and references, where higher values indicate a stronger positive correlation.
Lower values of RMSE, MAE, and MAPE demonstrate higher prediction accuracy. These metrics are defined as follows:
@@
To distinguish these models in subsequent experiments, we refer to the single-channel models as CNN-ConvLSTM (CNN-ConvLSTM-FCN), Conv-BiLSTM (Multi-scale Conv-BiLSTM), WaveTransNet, ViT-Wave, $\text{MLP}_{s}$ and $\text{CNN}_{s}$. The four-channel models are denoted as $\text{MLP}_{f}$, $\text{CNN}_{f}$ and SCAWaveNet. ${\rm CNN}_{s}$ and ${\rm MLP}_{s}$ are designed with the same number of layers and parameters settings as ${\rm CNN}_{f}$ and ${\rm MLP}_{f}$, respectively, to facilitate a fair comparison of different modeling approaches. Additionally, we employ two modeling strategies, i.e., Channel Independence (CI) and Channel Dependence (CD), within the four-channel models to explore how inter-channel information interaction influences model performance. The CI strategy extracts features from each channel independently, ignoring inter-channel correlations. The CD strategy treats all channels as an entity, allowing full interaction and dependency among channels during modeling.
@@
In the experiments, we used the PyTorch framework to train, validate, and test four-channel models. The optimizer is AdamW with a weight decay of $10^{-5}$. During training, batch size is set to 512, and the maximum number of epochs is 75. To avoid model overfitting and enhance computational efficiency, an early stopping strategy with a patience of 15 epochs is employed. The best model checkpoint is determined by the lowest average RMSE across the four channels. The Huber loss is adopted with a hyperparameter $\delta=2.0$. Learning rates varied between models due to architectural differences: $1.4\times 10^{-4}$ for SCAWaveNet and $10^{-4}$ for $\text{MLP}_{f}$ and $\text{CNN}_{f}$ (as their significant performance degrades when the learning rate is set to $1.4\times 10^{-4}$). For single-channel models, we maintain the same deep learning frameworks and experimental settings as those used in original studies to ensure methodological consistency. The four single-channel models are trained using data from each of the four channels. All models were developed on a single NVIDIA GeForce RTX 4080 GPU.
@@ V-B ERA5 Data Evaluation
Based on the ERA5 dataset, both comparative experiments and ablation experiments are provided in detail. To evaluate the algorithm’s performance, we also analyzed the number of parameters in the model. In addition, the analysis in this section does not consider the impact of environmental factors.
@@ V-B1 Quantitative Experiments
Figure 8: Density scatter plots on CYGNSS-ERA5 test set: (a) CNN-ConvLSTM, (b) ViT-Wave, (c) WaveTransNet, (d) $\text{MLP}_{f}$-CD, (e) $\text{CNN}_{f}$-CD, (f) SCAWaveNet-CI, (g) SCAWaveNet-CD, (h) SCAWaveNet-CD-wind.
@@
To analyze in more detail, we compare the prediction performance of selected models across varying SWH conditions. Based on the results in Table II, we select seven high-performing models for subsequent experiments. Specifically, CNN-ConvLSTM, ViT-Wave, and WaveTransNet are selected from the single-channel models, while $\text{MLP}_{f}$-CD, $\text{CNN}_{f}$-CD, SCAWaveNet-CI, and SCAWaveNet-CD are selected from the four-channel models. Fig. 7 presents histograms of the average RMSE and Bias values for the selected models across four channels within different ERA5 SWH ranges. RMSE and Bias exhibit distinct variation patterns: RMSE values remain relatively low in the 0-3 m range, but increase progressively when SWH >3 m. In contrast, models tend to overestimate SWH in the 0-2 m range, and this overestimation tendency weakens as the SWH increases. Once the SWH >2 m, underestimation begins to dominate and becomes increasingly pronounced at higher SWH values.
@@ V-B2 Qualitative Experiments
As shown in Fig. 8(g), the fitted line of SCAWaveNet-CD is closest to the ideal line compared to other models, particularly in the 0–1 m range, consistent with the findings in Fig. 7. While SCAWaveNet-CI (Fig. 8(f)) also performs well at low SWH, it tends to underestimate SWH when SWH >6 m. SCAWaveNet-CD exhibits fewer underestimations, suggesting that the incorporation of channel attention effectively alleviates the high-SWH underestimation. WaveTransNet (Fig. 8(c)) and $\text{CNN}_{f}$-CD (Fig. 8(e)) exhibit similar scatter distributions. This indicates that spatial attention in WaveTransNet contributes to effective feature extraction, and channel information interaction further enhances the representational capacity of CNNs. The fitted line of $\text{MLP}_{f}$-CD (Fig. 8(d)) deviates from the ideal line, which exhibits the most severe underestimation. This implies that MLPs capture highly localized features in spatial and channel dimensions. ViT-Wave (Fig. 8(b)), which extracts features only from DDMs without processing APs, exhibits poor performance, particularly at high SWH. In contrast, CNN-ConvLSTM (Fig. 8(a)), which uses MLP for APs feature extraction, has a fitted line closer to the ideal line than that of ViT-Wave. This further demonstrates the importance of APs feature extraction.
In terms of global retrieval errors, our algorithm also demonstrates superiority, as shown in Fig. 9. Fig. 9(a) illustrates the global distribution of the ERA5 test set. ERA5 SWH exhibits a hemispheric differences: lower values are generally observed in the Northern Hemisphere, whereas higher values prevail in the Southern Hemisphere, and increase with latitude. Additionally, the northwestern Indian Ocean in the Northern Hemisphere has relatively higher SWH due to the influence of monsoonal currents. Figs. 9(b)-(i) present the global bias distribution between model predictions and ERA5 values. To further examine the spatial patterns of the bias, four key regions are selected for detailed analysis: two with potential overestimation (Regions 1 and 2) and two with underestimation (Regions 3 and 4). Region 1 covers the Gulf of Mexico and the northwestern Caribbean Sea, Region 2 includes the South China Sea, the Philippine Sea, and the Coral Sea, Region 3 is located in the central South Pacific Ocean and Region 4 lies in the central-eastern Indian Ocean.
@@
DDMs APs RMSE (m) MAE (m) Bias (m) CC MAPE (%)
@@ V-B3 Ablation Experiments
To demonstrate the effectiveness of each component in SCAWaveNet, Table III compares the performance of SCAWaveNet-CD across different configurations. In the table, “DDMs” refers to the DDM branch, and“APs” refers to the APs branch. Within the “APs” category, “M”, “R” and “G” correspond to DDM-related, Receiver-related, and Geometry-related APs, respectively. The results show that the model integrating both branches outperforms single-branch configurations, confirming the complementary roles of these branches in enhancing model performance. Specifically, when using only the DDM branch, the model achieves an average RMSE of 0.526 m; in contrast, incorporating the APs branch reduces this RMSE to 0.452 m, indicating that the APs branch contributes more significantly to performance improvement. Further analysis of the APs branch reveals that DDM-related APs yield the most substantial contribution to performance enhancement, followed by Geometry-related APs and finally Receiver-related APs. Moreover, combining multiple types of APs leads to additional performance gains, demonstrating the benefit of incorporating all three types of APs into the modeling process.
@@ V-B4 Model Parameters and Complexity Analysis
In addition to quantitative and qualitative analyses of model performance, we further investigate model parameters and computational complexity. As shown in Table IV, “Params” denotes the total number of model parameters, including all weights and biases in the network, and “FLOPs” denotes the number of floating-point operations required for a single forward pass, which quantifies the computational complexity of the model. Among four-channel models, SCAWaveNet achieves the smallest parameter count and lowest FLOPs. Although SCAWaveNet-CD has approximately double the parameter count of SCAWaveNet-CI due to channel-wise interactions, their FLOPs are comparable, indicating that SCAWaveNet-CD maintains favorable computational efficiency. For single-channel models, WaveTransNet yields the lowest parameter count and FLOPs. However, SCAWaveNet-CI outperforms WaveTransNet in both parameter count and complexity, while SCAWaveNet-CD achieves comparable FLOPs but superior overall performance. These results underscore the advantage of SCAWaveNet in balancing model performance, parameter count, and computational complexity.

### v2

@@ III-B Lightweight Attention-based Auxiliary Parameters Branch
Then, $A_{e}$ is split into two equal-sized vectors, denoted as $A_{1}$ and $A_{2}$, which are fed into two distinct attention blocks. Each block includes two projection layers followed by a sigmoid activation function, designed to extract attention weights along the spatial and channel dimensions. Specifically, $A_{1}$ undergoes an up-projection to a higher-dimensional space first, then is down-projected back to its original dimension, and finally passes through a sigmoid function to generate spatial attention weights. For $A_{2}$, it is first transposed to feed into the attention block, and then transposed back to its original shape after extracting the attention weights. This process is defined as:
@@
where $\otimes$ denotes element-wise multiplication. The output $A^{\prime}$ serves as the enhanced APs feature, which integrates both spatial and channel-level characteristics.
@@ III-C Feature Fusion and Task Head
In the previous branches, features from the DDMs and APs were independently extracted to capture spatial and channel-wise information. These two types of features are concatenated along the channel dimension to form a unified representation. This feature representation is then fed into the task head, which is an MLP with 9 hidden layers. The output includes SWH predictions across four channels, expressed as:
@@
where $y$ and $\hat{y}$ denote the ground truth and the predicted values, respectively, $\delta$ is a hyperparameter that controls the transition between the Mean Squared Error (MSE) and Mean Absolute Error (MAE) characteristics. The Huber loss integrates the strengths of MSE and MAE: it behaves quadratically for small for small errors (ensuring smooth gradients) and linearly for large errors (reducing sensitivity to outliers). By tuning $\delta$, the loss function can balance the emphasis on large and small errors. In this study, the training objective is to optimize SWH predictions across all channels. Therefore, the loss is computed using the $\hat{y}$ and $y$ values of the four channels.
@@ IV-A Data Sources
In this study, GNSS-R observations are obtained from the CYGNSS Level 1 Version 3.2 science data product, provided by the Physical Oceanography Distributed Active Archive Center. CYGNSS, a NASA Earth System Science Pathfinder mission, consists of a constellation of eight microsatellites. These satellites provide near-continuous global coverage through an orbital inclination of approximately $35^{\circ}$ relative to the equator, with an average revisit time of 7 hours. This inclination enables CYGNSS to measure ocean parameters between nearly $38^{\circ}$ and $38^{\circ}$ latitude [49].
One of the reference datasets used in this study is the ERA5 dataset, the fifth-generation global climate and weather reanalysis product developed by the European Centre for Medium-Range Weather Forecasts[50]. ERA5 provides three types of SWH products: the combined SWH of wind waves and swell, the SWH of wind waves alone, and the SWH of swell alone. In this study, the combined SWH dataset are employed, which offer a spatial resolution of $0.5^{\circ}\times 0.5^{\circ}$ and an hourly temporal resolution. These datasets serve as the foundation for model training, validation, and testing.
In parallel, buoy data obtained from the National Data Buoy Center (NDBC) are used as real-world observations. The NDBC operates 1,350 stations, of which 1,056 have reported in the past 8 hours. We specifically utilized standard meteorological NDBC data, with each station reporting SWH measurements at 30- or 60-minute intervals. Due to the limited availability of buoy data, these measurements are used exclusively as a test set to evaluate the real-world performance of the SWH retrieval model.
@@ IV-B Data Preprocessing
After acquiring CYGNSS, ERA5 and NDBC data, a three-step data preprocessing procedure is applied: (1) quality control; (2) channel alignment; (3) CYGNSS-ERA5 and CYGNSS-Buoy data matching.
@@ IV-B1 Quality Control
Figure 5: Data processing workflow used in this study.
@@
  • Remove observations where any of the first 28 bits in the quality control flag equals to 1.
@@ IV-B2 Channel Alignment
  • If data from any channel are missing at a certain timestamp, the entire observation for that timestamp is discarded.
@@ IV-B3 Data Matching
  • CYGNSS-Buoy data matching: The temporal and spatial matching thresholds for CYGNSS and NDBC buoy data are set to 30 minutes and 25 km, whereby the CYGNSS specular point must be within 25 km of the buoy location and the observation time must fall within ± 30 minutes of the buoy measurement. Due to the stringent requirement of aligning all four CYGNSS channels with identical timestamps, data matching is performed independently for each channel.
Due to the fact that the CYGNSS data exhibit signal saturation when SWH >8 m, these measurements offer limited modeling value. Therefore, we exclude all data points with SWH values above this threshold.
@@ IV-C Data Split
In this study, the temporal coverage of CYGNSS and ERA5 data spans from August 1, 2019, to July 31, 2022. To ensure fairness in model training and to account for potential seasonal variations, the training, validation, and test sets were strictly divided into non-overlapping time periods. Specifically:
@@
  • The test set ranges from August 1, 2021, to July 31, 2022.
Given that the dataset contains four-channel data, training a model on the full yearly dataset would be computationally expensive. Therefore, we randomly sampled 3 million, 0.5 million, and 3 million data points to form the final training, validation, and test sets, respectively. Furthermore, to ensure consistency in model evaluation, the buoy data was aligned with the temporal range of the test set.
@@ V-A Experimental Setup
  • CC quantifies the linear relationship between predictions and references, with higher values indicating a stronger positive correlation.
Lower values of RMSE, MAE, and MAPE indicate a higher prediction accuracy. These metrics are defined as follows:
@@
To distinguish these models in subsequent experiments, we refer to the single-channel models as CNN-ConvLSTM (CNN-ConvLSTM-FCN), Conv-BiLSTM (Multi-scale Conv-BiLSTM), WaveTransNet, ViT-Wave, $\text{MLP}_{s}$ and $\text{CNN}_{s}$. The four-channel models are denoted as $\text{MLP}_{f}$, $\text{CNN}_{f}$ and SCAWaveNet. ${\rm CNN}_{s}$ and ${\rm MLP}_{s}$ are designed with the same number of layers and parameters settings as ${\rm CNN}_{f}$ and ${\rm MLP}_{f}$, respectively, to facilitate a fair comparison of different modeling approaches. Additionally, we employ two modeling strategies, i.e., Channel Independence (CI) and Channel Dependence (CD), within the four-channel models to explore how inter-channel information interaction influences model performance. The CI strategy extracts features from each channel independently, ignoring inter-channel correlations. The CD strategy treats all channels as a unified entity, allowing full interaction and dependency among channels during modeling.
@@
In the experiments, we employed the PyTorch framework to train, validate, and test four-channel models. The optimizer is AdamW with a weight decay of $10^{-5}$. During training, the batch size is set to 512, and the maximum number of epochs is 75. To avoid model overfitting and enhance computational efficiency, an early stopping strategy with patience of 15 epochs is applied. The best model checkpoint is determined by the lowest average RMSE across four channels. The Huber loss is adopted with a hyperparameter $\delta=2.0$. Learning rates were adjusted according to the model architectures: $1.4\times 10^{-4}$ for SCAWaveNet and $10^{-4}$ for $\text{MLP}_{f}$ and $\text{CNN}_{f}$ (as their significant performance degrades when the learning rate is set to $1.4\times 10^{-4}$). For single-channel models, we maintain the same deep learning frameworks and experimental settings as those used in the original studies to ensure methodological consistency. The four single-channel models were trained separately using data from each of the four channels. All models were developed on a single NVIDIA GeForce RTX 4080 GPU.
@@ V-B ERA5 Data Evaluation
Based on the ERA5 dataset, both comparative experiments and ablation experiments are provided in detail. To evaluate the algorithm’s performance, we also analyzed the number of parameters in the model.
@@ V-B1 Quantitative Experiments
Figure 8: Density scatter plots of model predictions for the CYGNSS-ERA5 channel 1 test set: (a) CNN-ConvLSTM, (b) ViT-Wave, (c) WaveTransNet, (d) $\text{MLP}_{f}$-CD, (e) $\text{CNN}_{f}$-CD, (f) SCAWaveNet-CI, (g) SCAWaveNet-CD, (h) SCAWaveNet-CD-wind.
@@
To analyze in more detail, we compare the prediction performance of selected models across varying SWH conditions. Based on the results in Table II, we select seven high-performing models for subsequent experiments. Specifically, CNN-ConvLSTM, ViT-Wave, and WaveTransNet are selected from the single-channel models, while $\text{MLP}_{f}$-CD, $\text{CNN}_{f}$-CD, SCAWaveNet-CI, and SCAWaveNet-CD are selected from the four-channel models. Fig. 7 presents histograms of the average RMSE and Bias values for the selected models across four channels within different ERA5 SWH ranges. RMSE and Bias exhibit distinct variation patterns: RMSE values remain relatively low within 0-3 m, but increase progressively when SWH >3 m. In contrast, models tend to overestimate SWH in the 0-2 m range, and this overestimation tendency weakens as the SWH increases. Once the SWH >2 m, underestimation begins to dominate and becomes increasingly pronounced at higher SWH values.
@@ V-B2 Qualitative Experiments
As shown in Fig. 8(g), the fitted line of SCAWaveNet-CD is closest to the ideal line compared to other models, especially in the 0–1 m range, consistent with the findings in Fig. 7. While SCAWaveNet-CI (Fig. 8(f)) also performs well at low SWH, it tends to underestimate SWH when SWH >6 m. SCAWaveNet-CD exhibits fewer underestimations, suggesting that the incorporation of channel attention effectively alleviates the high-SWH underestimation. WaveTransNet (Fig. 8(c)) and $\text{CNN}_{f}$-CD (Fig. 8(e)) exhibit similar scatter distributions. This indicates that spatial attention in WaveTransNet contributes to effective feature extraction, and channel information interaction further enhances the representational capacity of CNNs. The fitted line of $\text{MLP}_{f}$-CD (Fig. 8(d)) deviates from the ideal line, which exhibits the most severe underestimation. This implies that MLPs capture highly localized features in spatial and channel dimensions. ViT-Wave (Fig. 8(b)), which extracts features only from DDMs without processing APs, exhibits poor performance, particularly at high SWH. In contrast, CNN-ConvLSTM (Fig. 8(a)), which uses MLP for APs feature extraction, has a fitted line closer to the ideal line than that of ViT-Wave. This further demonstrates the importance of APs feature extraction.
In terms of global retrieval errors, our algorithm also demonstrates superiority, as shown in Fig. 9. Fig. 9(a) illustrates the global distribution of the ERA5 test set. ERA5 SWH exhibits a hemispheric difference: lower values are generally observed in the Northern Hemisphere, whereas higher values prevail in the Southern Hemisphere, and increase with latitude. Additionally, the northwestern Indian Ocean in the Northern Hemisphere has relatively higher SWH values due to the influence of monsoonal currents. Figs. 9(b)-(i) present the global bias distribution between model predictions and ERA5 values. To further examine the spatial patterns of the bias, four key regions are selected for detailed analysis: two with potential overestimation (Regions 1 and 2) and two with underestimation (Regions 3 and 4). Region 1 covers the Gulf of Mexico and the northwestern Caribbean Sea, Region 2 includes the South China Sea, the Philippine Sea, and the Coral Sea, Region 3 is located in the central South Pacific Ocean and Region 4 lies in the central-eastern Indian Ocean.
@@
DDM AP RMSE (m) MAE (m) Bias (m) CC MAPE (%)
@@ V-B3 Ablation Experiments
To demonstrate the effectiveness of each component in SCAWaveNet, Table III compares the performance of SCAWaveNet-CD across different configurations. In the table, “DDM” refers to the DDM branch and“AP” refers to the AP branch. Within the “AP” category, “M”, “R” and “G” correspond to DDM-related, Receiver-related, and Geometry-related APs, respectively. The results indicate that the model integrating both branches outperforms single-branch configurations, confirming the complementary roles of these branches in enhancing model performance. Specifically, when using only the DDM branch, the model achieves an average RMSE of 0.526 m; in contrast, incorporating the AP branch reduces this RMSE to 0.452 m, indicating that the AP branch contributes more significantly to performance improvement. Further analysis of the AP branch reveals that DDM-related APs yield the most substantial contribution to performance enhancement, followed by Geometry-related APs and finally Receiver-related APs. Moreover, combining multiple types of AP leads to additional performance gains, demonstrating the benefit of incorporating all three types of AP into the modeling process.
@@ V-B4 Model Parameters and Complexity Analysis
In addition to quantitative and qualitative analyses of model performance, we further investigate model parameters and computational complexity. As shown in Table IV, “Params” denotes the total number of model parameters, including all weights and biases in the network, and “FLOPs” denotes the number of floating-point operations required for a single forward pass, which quantifies the computational complexity of the model. Among four-channel models, SCAWaveNet achieves the smallest parameter count and the lowest FLOPs. Although SCAWaveNet-CD has approximately double the parameter count of SCAWaveNet-CI due to channel-wise interactions, their FLOPs remain comparable, indicating that SCAWaveNet-CD maintains favorable computational efficiency. For single-channel models, WaveTransNet yields the lowest parameter count and FLOPs. However, SCAWaveNet-CI outperforms WaveTransNet in both parameter count and complexity, while SCAWaveNet-CD achieves comparable FLOPs but superior overall performance. These results underscore the advantage of SCAWaveNet in balancing model performance, parameter count, and computational complexity.